# Generative AI for Deep Reinforcement Learning: Framework, Analysis, and Use Cases

### v1

@@ Abstract
Generative AI, Deep reinforcement learning, near-filed communication, optimization.
@@ I Introduction
  • Generalization Ability: GAI can effectively apply the knowledge and representations learned from one task to other tasks, which accelerates the learning process and improves the performance in new environments.
@@
Motivated by this, we attempt to propose a comprehensive tutorial. To the best of our knowledge, this is the first work to systematically demonstrate the use of GAI to enhance DRL. The contributions can be summarized as follows:
  • We present the principles of important and widely used GAI models and DRL algorithms. Then we briefly describe several applications of DRL enhanced with GAI.
  • We explore different integrations of GAI in DRL, focusing on how DRL can be enhanced from a data and strategy perspective.
  • We propose a noval framework for DRL enhancement by GAI. In addition, we construct a case study to demonstrate the effectiveness of the proposed framework.
Figure 1: Summary of typical GAI models.
@@ II-A GAI Models
DRL Algorithm Analysis Applications boosted by GAI
@@
DQN Off-policy Estimating Q-function using DNN and handle discrete action space $\checkmark$ Simple to implement $\checkmark$ Good generality $\times$× Training instability Signal Processing: GAI is utilized to extract crucial information carried by multimodal signal data. Wireless Communications: GAI can generate data to reduce the overhead of interacting with the real world. Edge-based AIGC Service: GAI can be used to generate scheduling policy to improve the effectiveness of AIGC service. Sensor Networks: GAI is utilized to improve the decision-making ability of DRL in cognitive internet of things network.
@@ II-C Applications of GAI-enhanced DRL
Figure 2: Summary of data and policy performance improved by GAI in DRL. GAI can be used to improve the data performance of DRL through data augmentation, handling variable state spaces, extracting features, and simulating environments. Moreover, GAI can enhance the policy performance of DRL by improving the policy network, learning from multimodal data, handling hybrid actions, and transferring knowledge across different tasks.
@@ III-B2 Multimodal Learning
SAC $\bullet$ Extract latent features. $\bullet$ Encode historical information and hybrid space. $\bullet$ Apply GDM structure to policy network [3]. $\bullet$ Simulate expert behavior to generate optimal policy. $\bullet$ Provide a simulated environment for SAC agent. $\bullet$ Map high-dimensional states into the latent vectors [13]. $\bullet$ Encode tasks and infer tasks.
@@ III-B4 Transfer Learning
Figure 3: Proposed Framework of GAI-enhanced DRL on UAV-assisted integrated near-field/far-field communication. The generators of GAN are used as critic and target critic network. VAE is used to reduce the dimensionality of high-dimensional input and process the hybrid action space. The attention mechanism of Transformer is used to enhance the actor network. The reverse process of GDM is used to replace the actor network.
@@ III-C Discussion
  • Improved Decision Ability: The reverse process of the diffusion model is well-suited to the policy networks of DRL, where multi-round iterations and denoising processes help agents explore better actions. Additionally, GAI has inspired the implementation of transfer learning in DRL to facilitate interactions across similar domains
IV Case Study: GAI-enhanced DRL in Near-filed Communication
@@ IV-A2 VAE-enhanced DRL
We use VAE to reduce the dimensionality of the high-dimensional state space to reduce the computational complexity issue in DRL. In this case, we train the VAE with data and use the decoder to extract representations of the state space, which are then used as inputs for the actor and critic networks. Additionally, VAE can construct a latent representation space for continuous parameters conditioned on state and embedding of discrete actions to handle hybrid actions. This method enables more efficient policy learning and action selection in complex high-dimensional environments.
@@ IV-A3 Transformer-enhanced DRL
We enhance the actor network of DRL by using Transformer. Specifically, we replace the Multi-Layer Perceptron (MLP) with a network based on the attention mechanism of Transformer to analyze the current state in the environment. This approach helps to capture important information more efficiently by dynamically focusing on relevant parts of the input data.
@@ IV-A4 GDM-enhanced DRL
We improve the policy network of DRL by employing the reverse process of GDM. Specifically, we treat the policy network as a denoiser, progressively adding denoising noise to the initial Gaussian noise to recover or discover the optimal actions. GDM effectively captures the dependencies between state and action spaces, aiding the agent in selecting appropriate actions based on the observed states.
Figure 4: The convergence performance of case study. The convergence curves of GAI-enhanced DRL methods in continuous action space for reward, transmission rate and power consumption are shown in (a), (b) and (c), respectively. The convergence curves of reward, transmission rate and power consumption in discrete space are presented in (d), (e) and (f), respectively. The convergence curves for reward, transmission rate, and power consumption in discrete space are displayed in (g), (h), and (i), respectively.
@@ IV-C Result Analysis
To investigate an impact of different types of action spaces, we constructed three experiments to explore the performance of GAI-enhanced DRL in continuous action space (i.e., continuous UAV trajectory design and power allocation), discrete action space (i.e., way-point discrete UAV trajectory design and power allocation), and hybrid action spaces (i.e., discrete UAV trajectory design and continuous power allocation).
IV-C1 Continuous Action Space
@@
Figs. 4 (a), (b) and (c) show the convergence curves of reward, transmission rate and power consumption in continuous action space, respectively. As an improved version of DDPG, TD3 significantly enhances policy stability and performance through the use of twin Q-networks and delayed policy updates. Therefore, in our case study, we select TD3 as the foundational DRL algorithm. We can observe that the GAN-enhanced TD3 algorithm shows significant improvements over the original TD3 algorithm, and this improvement is due to the novel adversarial training of the critic and target critic networks, which enhances the process of state-action value estimation. However, it is unable to effectively learn the constraints on power within the continuous solution space, resulting in penalties that cause fluctuations in the reward. Furthermore, we observe that VAE improves the performance of the TD3 algorithm, which is attributed to the capability of VAE to reduce the dimensionality of the high-dimensional state space. However, the process of transforming the high-dimensional state space into a low-dimensional latent space results in instability. Additionally, we find that the Transformer-enhanced TD3 algorithm barely improves the performance of the original TD3 algorithm. This may be because Transformer is better suited for handling long sequence problems, which may not align well with the current optimization problem. In addition, we observe that GDM-enhanced-TD3 achieves the best reward values and transmission rates compared to other algorithms. This is because GDM can accurately capture the underlying data distribution, providing a more effective representation of the environment. Moreover, the unique structure of GDM, which involves a diffusion process, offers a more stable and efficient learning process.
IV-C2 Discrete Action Space
@@
In this case, the convergence curves of reward, transmission rate and power consumption in the discrete action space are shown in Figs. 4 (d), (e) and (f), respectively. Unlike its performance in continuous action space, GAN-enhanced GAI demonstrates superior convergence speed and quality in all of metrics above. This is because the discrete action space we designed inherently satisfies the power constraints of the optimization problem, eliminating the need for GAN to learn these constraints. However, this also indirectly highlights the weakness of GAN in learning constraint conditions directly related to optimization variables. Additionally, the convergence curves of VAE-enhanced GAI in discrete space are relatively unstable. This could be due to the fact that in discrete action spaces, the decision boundaries are usually more complex and require precise state information to make correct action choices. However, the latent space obtained after dimensionality reduction by the VAE model may not provide sufficiently accurate information to correctly distinguish different discrete actions, leading to decision biases. Moreover, we observe from the figure that the GDM-based algorithm still achieves the excellent performance, which indicates that it can handle discrete actions well.
IV-C3 Hybrid Action Space
@@
Figs. 4 (g), (h) and (i) show the convergence curves of reward, transmission rate and power consumption in hybrid action space, respectively. We can observe that the GDM-enhanced TD3 is optimal in terms of the convergence speed and stability on the hybrid action space, and it converges to the best values with regard to reward, transmission rate, and power consumption. In the hybrid action space, the discrete and continuous actions are not independent, and different discrete actions correspond to different optimal continuous actions [12]. Therefore, to make the best hybrid action decisions, it is essential to explore the latent relationships between discrete and continuous actions. Evidently, the strong capability of GDM in learning the underlying data relationships enhances the accuracy and stability of TD3 decisions, enabling GDM-enhanced TD3 to achieve outstanding performance in the hybrid action space. The comparison above is limited due to the space constraint. More extensive investigation would be needed from other perspectives and system settings.
@@ V-A Efficiency Enhancement of GAI-enhanced DRL
Although improving performance significantly, the introduction of GAI models inevitably increases the computational complexity of the original DRL algorithms, which clearly leads to a need for more computational resources and longer training times. Therefore, enhancing the computational efficiency of GAI-enhanced DRL algorithms is crucial. In this case, utilizing GAI to generate more efficient exploration strategies to reduce the number of training samples might be a feasible approach.
@@ V-B GAI-assisted Reward Design
As a key component of DRL, the quality of reward function design significantly impacts the performance of DRL. However, reward functions are usually manually set, and finding an appropriate reward function often requires multiple trials. Therefore, GAI can be used to design complex reward functions, improving accuracy and reducing the workload of manual design. Alternatively, inverse reinforcement learning can be considered to define the reward functions.
@@ V-C GAI-improved Transfer Learning
If the network architecture of DRL changes, we would have to retrain DRL from scratch, which results in a waste of computational resources. It is worth noting that GAI can apply knowledge learned from one task to another related task, thereby accelerating the learning process or improving performance. Therefore, leveraging GAI for knowledge transfer in DRL is a very promising direction.
@@ VI Conclusion
In this paper, we systematically introduced how GAI can be applied to DRL to improve the performance. Specifically, we first introduced the basic principles of several GAI models and DRL algorithms. Subsequently, we discussed how GAI enhances DRL algorithms from both the data and policy perspectives. Then, we proposed a framework to explore the improvement of DRL algorithm by four typical GAI models, namely GAN, GDM, VAE, and Transformer. Moreover, we constructed a case study on near-field communication to validate the effectiveness of GAI-enhanced DRL. Experimental results demonstrated that GDM shows the most significant enhancement for DRL. Finally, we introduced three future directions to advance the wider application of GAI in DRL.

### v2

@@ Abstract
Generative AI, Deep reinforcement learning, near-field communication, optimization.
@@ I Introduction
  • Generalization Ability: GAI can efficiently transfer the knowledge and representations acquired from one task to others, thereby accelerating the learning process and enhancing performance in new environments.
Figure 1: Summary of typical GAI models.
@@
Motivated by this, we attempt to propose a comprehensive tutorial. To the best of our knowledge, this is the first work to comprehensively summarize how DRL algorithms can be enhanced using various GAI models. The contributions can be summarized as follows:
  • We present the principles of important and widely used GAI models and DRL algorithms. Then, we briefly describe several applications of DRL enhanced with GAI.
  • We explore different integration of GAI in DRL, focusing on how DRL can be enhanced from a data and strategy perspective. Moreover, we comprehensively summarize the advantages and disadvantages of the GAI-enhanced DRL algorithm.
  • We propose a novel framework which deeply demonstrates the integration process of several common GAI models individually in the DRL algorithm. Moreover, we show the role of integrating multiple GAI models simultaneously in DRL to solve different challenges in DRL algorithms. Furthermore, we construct a case study to demonstrate the effectiveness of the proposed framework.
@@ II-A GAI Models
DRL Algorithm Analysis
@@
DQN Off-policy Estimating Q-function using DNN and handle discrete action space $\checkmark$ Simple to implement $\checkmark$ Good generality $\times$× Training instability
@@ II-C Applications of GAI-enhanced DRL
Figure 2: Summary of data and policy performance improved by GAI in DRL.
@@ III-B2 Multimodal Learning
SAC $\bullet$ Extract latent features. $\bullet$ Encode historical information and hybrid space. $\bullet$ Apply GDM structure to policy network [3]. $\bullet$ Simulate expert behavior to generate optimal policy. $\bullet$ Provide a simulated environment for SAC agent [9]. $\bullet$ Map high-dimensional states into the latent vectors. $\bullet$ Encode tasks and infer tasks.
@@ III-B4 Transfer Learning
Figure 3: Proposed Framework of GAI-enhanced DRL on UAV-assisted integrated near-field/far-field communication.
@@ III-C Discussion
  • Improved Decision Ability: The reverse process of the diffusion model is well-suited to the policy networks of DRL, where multi-round iterations and denoising processes help agents explore better actions. Additionally, GAI has inspired the implementation of transfer learning in DRL to facilitate interactions across similar domains.
Although incorporating GAI into DRL addresses some issues faced by DRL, it also introduces some disadvantages as follows:
  • High Computation Complexity: Integrating GAI models into DRL algorithms may introduce additional computational costs, such as tuning the network architecture of GAI models and a large number of hyperparameters, which can increase computational overhead during training. Moreover, the training process of GAI models requires iterative optimization, which leads to an increase in the overall training time of DRL.
  • High Resource Demand: GAI models typically require substantial memory to store model weights, parameters, and intermediate results, which is particularly pronounced in data-intensive applications. Moreover, GAI models often have complex structures and require high-performance computing resources to support their training when performing computationally intensive tasks.
Therefore, GAI-enhanced DRL is suitable for solving the issues of low sample efficiency and insufficient exploration by generating diverse training data and high-quality actions with sufficient computational resources.
IV Case Study: GAI-enhanced DRL in Near-field Communication
@@ IV-A2 VAE-enhanced DRL
We use VAE to reduce the dimensionality of the high-dimensional state space in DRL. Specifically, we input the high-dimensional state data into the encoder of the VAE to extract a lower-dimensional latent variable. Subsequently, the decoder attempts to reconstruct the original state data from the latent variable. By minimizing the discrepancy between the reconstructed data and the original data, the VAE is able to maintain the quality of the lower-dimensional latent variable while performing dimensionality reduction. Additionally, VAE can construct a latent representation space for continuous parameters conditioned on state and embedding of discrete actions to handle hybrid actions.
@@ IV-A3 Transformer-enhanced DRL
We enhance the actor network of DRL by using Transformer. Specifically, we replace the MLP with a network based on the attention mechanism of Transformer to analyze the current state in the environment. This enables Transformer-enhanced DRL to more effectively capture long-range dependencies within sequences, making it well-suited for decision-making tasks that require long-term memory or complex historical information.
@@ IV-A4 GDM-enhanced DRL
We improve the policy network of DRL by employing the reverse process of GDM. Specifically, we consider the policy network as a denoiser, progressively adding denoising noise to the initial Gaussian noise to recover or discover the optimal actions. Benefiting from the generative ability and expressiveness of GDM, it can generate diverse and high-quality actions, which makes it suitable for improving DRL algorithms with limited exploration capabilities.
Figure 4: The convergence curves of rewards, rates, and power consumption, as well as the training time, of the case study under three different action spaces. Specifically, the results in continuous action space are shown in (a), (b), (c) and (d), respectively. The results in discrete action space are shown in (e), (f), (g) and (h), respectively. The results in hybrid action space are shown in (i), (j), (k) and (l), respectively.
@@ IV-C1 Analysis of Different Types of Action Spaces
We explore the performance of GAI-enhanced DRL in continuous action space (i.e., continuous UAV trajectory design and power allocation), discrete action space (i.e., way-point discrete UAV trajectory design and power allocation), and hybrid action spaces (i.e., discrete UAV trajectory design and continuous power allocation).
@@
Continuous Action Space: Figs. 4(a), (b) and (c) show the convergence curves of reward, transmission rate and power consumption in continuous action space, respectively. As an improved version of DDPG, TD3 significantly enhances policy stability and performance through the use of twin Q-networks and delayed policy updates. Therefore, in our case study, we select TD3 as the foundational DRL algorithm. We can observe that the GAN-enhanced TD3 algorithm shows significant improvements over the original TD3 algorithm, and this improvement is due to the novel adversarial training of the critic and target critic networks, which enhances the process of state-action value estimation. However, it is unable to effectively learn the constraints on power within the continuous solution space, resulting in penalties that cause fluctuations in the reward. Moreover, we observe that VAE improves the performance of the TD3 algorithm, which is attributed to the capability of VAE to reduce the dimensionality of the high-dimensional state space to extract critical information. However, the process of transforming the high-dimensional state space into a low-dimensional latent space inevitably results in loss of information and leads to instability.. Furthermore, we find that the Transformer-enhanced TD3 algorithm barely improves the performance of the original TD3 algorithm. This may be because Transformer is better suited for handling long-range problems, which may not align well with the current optimization problem. In addition, we observe that GDM-enhanced-TD3 achieves the best reward values and transmission rates compared to other algorithms. This is because GDM can accurately capture the underlying data distribution, providing a more effective representation of the environment. Moreover, the unique structure of GDM, which involves a diffusion process, offers a more stable and efficient learning process.
@@
Discrete Action Space: The convergence curves of reward, transmission rate and power consumption in the discrete action space are shown in Figs. 4(e), (f) and (g), respectively. Unlike its performance in continuous action space, GAN-enhanced GAI demonstrates superior convergence speed and quality in all of metrics above. This is because the discrete action space we designed inherently satisfies the power constraints of the optimization problem, eliminating the need for GAN to learn these constraints. However, this also indirectly highlights the weakness of GAN in learning constraint conditions directly related to optimization variables. Moreover, the convergence curves of VAE-enhanced GAI in discrete space are relatively unstable. This could be due to the fact that in discrete action spaces, the decision boundaries are usually more complex and require precise state information to make correct action choices. However, the latent space obtained after dimensionality reduction by the VAE model may not provide sufficiently accurate information to correctly distinguish different discrete actions, leading to decision biases. Furthermore, we observe from the figure that the GDM-based algorithm still achieves the excellent performance, which indicates that it can handle discrete actions well.
@@
Hybrid Action Space: Figs. 4(i), (j) and (k) show the convergence curves of reward, transmission rate and power consumption in hybrid action space, respectively. We can observe that the GDM-enhanced TD3 is optimal in terms of the convergence speed and stability on the hybrid action space, and it converges to the best values with regard to reward, transmission rate, and power consumption. In the hybrid action space, the discrete and continuous actions are not independent, and different discrete actions correspond to different optimal continuous actions [12]. Therefore, to make the best hybrid action decisions, it is essential to explore the latent relationships between discrete and continuous actions. Evidently, the strong capability of GDM in learning the underlying data relationships enhances the accuracy and stability of TD3 decisions, enabling GDM-enhanced TD3 to achieve outstanding performance in the hybrid action space.
IV-C2 Analysis of Multiple GAI Models Combination
We further combine multiple GAI models in TD3 to explore their performance on the abovementioned three action spaces. Specifically, we propose the GDM-GAN-enhanced TD3 algorithm, which introduces GAN into the GDM-enhanced TD3 algorithm to improve the critic network. Based on this, we design the GDM-GAN-Transformer-enhanced TD3 algorithm, which improves the denoising network of GDM by using Transformer. Finally, the GDM-GAN-Transformer-VAE-enhanced TD3 algorithm is proposed, which introduces VAE into the GDM-GAN-Transformer-enhanced TD3 to reduce the dimensionality of the state space. The results are shown in Fig. 4 and corresponding results analysis are as follows:
  • Results of GDM-GAN-enhanced TD3 algorithm: As can be seen, the GDM-GAN-enhanced TD3 algorithm slightly outperforms the GDM-enhanced TD3 algorithm. This is because the GAN further enhances the quality of the estimation of state-action values by the critic network, which is beneficial for the agent to explore more effective actions.
  • Results of GDM-GAN-Transformer-enhanced TD3 algorithm: Introducing the Transformer does not lead to significant changes in the performance of the GDM-GAN-enhanced TD3 algorithm. This is because Transformer is more suitable for addressing long-range optimization problems and it does not provide substantial performance gains in our considered case study.
  • Results of GDM-GAN-Transformer-VAE-enabled TD3 algorithm: Adopting VAE results in the performance decline of the GDM-GAN-Transformer-enhanced TD3 algorithm. This decline results from the information loss caused by the dimensionality reduction of state data using VAE, which negatively impacts the GAN-based critic network that relies on detailed state information, leading to an overall performance decrease.
  • Training Time: The training time is also newly recorded in Fig. 4, and it can be seen from the figure that the introduced additional improved factors, i.e., the GAI models, inevitably increases training time. Moreover, we can observe that the training time of GDM-enhanced TD3 is significantly higher than that of the original TD3 algorithm. This is because the denoising time step in GDM leads to an increase in the training time of the neural network, while this also leads to a notable performance improvement. Similarly, although the training time of GDM-GAN-enhanced TD3 is slightly higher than that of the GDM-enhanced TD3, the former achieves better results, which is particularly beneficial for scenarios that require high decision accuracy.
Ultimately, all the GAI models are able to improve the performance of the original TD3 algorithm for dealing with the optimization problem in the case study. Moreover, more integration combination of GAI models in DRL can be explored in the future work.
@@ V-A Efficiency Enhancement of GAI-enhanced DRL
Although GAI models can improve the performance of DRL algorithms significantly, these models inevitably increases the computational complexity of the original DRL algorithms, which requires longer training time. Therefore, it is crucial to enhance the computational efficiency of GAI-enhanced DRL algorithms. One approach is to decompose the complex task into multiple simpler sub-tasks and distribute these sub-tasks across multiple machines or processors during the training process. This approach enhances efficiency by focusing on smaller and more manageable issues.
@@ V-B GAI-assisted Reward Design
As a key component of DRL, the quality of reward function design significantly impacts the performance of DRL. However, reward functions are usually manually set, and finding an appropriate reward function often requires multiple trials. In this context, the LLM integrated with retrieval-augmented generation (RAG) technique can offer a novel approach to improve the reward function design process. Specifically, the RAG model retrieves relevant information from existing reward function databases and research papers about DRL. Subsequently, the retrieved information is combined with the LLM to adjust and fine-tune the reward function for specific tasks or environments, thereby enhancing the accuracy of reward design.
@@ V-C GAI-improved Transfer Learning
If the application scenarios of DRL change, we usually have to retrain the DRL models from scratch, which results in a waste of computational resources. Notably, GAI shows great potential to solve this issue. Specifically, GAI can transfer the learned knowledge across different domains by identifying and exploiting common underlying structures that unify the distribution of newly generated data to a given distribution. Therefore, leveraging GAI to identify meta-knowledge that can be transferred across various domains helps bridge the gaps among different application scenarios of DRL [10515203].
V-D Exploration of More GAI Models for DRL
Different GAI models possess distinct characteristics and focuses. Therefore, it is essential to explore the integration of more GAI models into DRL algorithms to address various challenges faced by different DRL algorithms and improve the decision-making process. Based on this, refining the criteria for GAI model selection can ensure that the most suitable GAI model can be quickly chosen to address specific application scenarios and challenges within DRL. Moreover, it is necessary to develop appropriate evaluation metrics to assess the impact of GAI models on the performance of DRL algorithms.
@@ VI Conclusion
In this paper, we have introduced how GAI can be applied to DRL to improve the performance. Specifically, we have first introduced the basic principles of several GAI models and DRL algorithms. Subsequently, we have discussed how GAI enhances DRL algorithms from both the data and policy perspectives. Then, we have proposed a framework to explore the improvement of DRL algorithm by four typical GAI models, namely GAN, GDM, VAE, and Transformer. Moreover, we have constructed a case study on near-field communication to validate the effectiveness of GAI-enhanced DRL. Experimental results have demonstrated that GDM shows the most significant enhancement for DRL. Finally, we have introduced four future directions to advance the wider application of GAI in DRL.
Biographies
Geng Sun (sungeng@jlu.edu.cn) received a B.S. degree in Communication Engineering from Dalian Polytechnic University, China, and the Ph.D. degree in Computer Science from Jilin University, China, in 2011 and 2018, respectively. He was a visiting researcher in the School of Electrical and Computer Engineering at Georgia Institute of Technology, USA. He is currently a Professor in the College of Computer Science and Technology at Jilin University, and his research interests include UAV Networks, collaborative beamforming, generative AI and optimizations.
Wenwen Xie (xieww22@mails.jlu.edu.cn) received the B.S. degree in Computer Science and Technology from Hefei University of Technology, Hefei, China, in 2022. She is currently working toward the MS degree in Computer Science and Technology at Jilin University, Changchun, China. Her research interests include UAV communications, IRS beamforming and deep reinforcement learning.
Dusit Niyato (dniyato@ntu.edu.sg) is currently a professor in the School of Computer Science and Engineering, Nanyang Technological University, Singapore. He received the B.Eng. degree from King Mongkut’s Institute of Technology Ladkrabang (KMITL), Thailand, in 1999, and the Ph.D. in electrical and computer engineering from the University of Manitoba, Canada, in 2008. His research interests are in the areas of Internet of Things (IoT), machine learning, and incentive mechanism design.
Fang Mei (meifang@jlu.edu.cn) received the M.Sc. and Ph.D. degrees in computer science from Jilin University, Changchun, China, in 2005 and 2010, respectively. She is currently an Associate Professor in the College of Computer Science and Technology at Jilin University. Her research interests include intelligent information processing, multi-access edge computing, vehicular networks, and antenna arrays.
Jiawen Kang (kavinkang@gdut.edu.cn) received the Ph.D. degree from the Guangdong University of Technology, China, in 2018. He has been a postdoc at Nanyang Technological University, Singapore, from 2018 to 2021. He is currently a full professor at Guangdong University of Technology, China. His research interests focus on blockchain, security, and privacy protection.
Hongyang Du (duhy@hku.hk) is an assistant professor at the Department of Electrical and Electronic Engineering, The University of Hong Kong. He received his Ph.D. degree from the College of Computing and Data Science, Energy Research Institute @ NTU, Nanyang Technological University, Singapore, in 2024. He serves as the Editor-in-Chief assistant of IEEE Communications Surveys & Tutorials (2022-2024), and the Guest Editor for IEEE Vehicular Technology Magazine. His research interests include edge intelligence, generative AI, and network management.
Shiwen Mao (smao@ieee.org) received his Ph.D. in electrical and computer engineering from Polytechnic University, Brooklyn, NY. He is a Professor and Earle C. Williams Eminent Scholar, and Director of the Wireless Engineering Research and Education Center at Auburn University. His research interests include wireless networks and multimedia communications.